# Improving Stochastic Cubic Newton with Momentum

### v1

@@ 1.1 Motivation
SCN with MVR–HB momentum
(ours) see Appendix F
$F_{0}\cdot\mathcal{O}\Bigl{(}{\color[rgb]{0.15234375,0.51171875,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.15234375,0.51171875,0.26171875}\bm{\frac{L^{1/2}}{\varepsilon^{3/2}}}}+{\color[rgb]{0.15234375,0.51171875,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.15234375,0.51171875,0.26171875}\bm{\frac{{\color[rgb]{0.75390625,0.18359375,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0.75390625,0.18359375,0.09765625}\bm{L_{g}}}+\sigma_{h}}{\varepsilon^{2}}}}+{\color[rgb]{0.15234375,0.51171875,0.26171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.15234375,0.51171875,0.26171875}\bm{\frac{{\color[rgb]{0.75390625,0.18359375,0.09765625}\definecolor[named]{pgfstrokecolor}{rgb}{0.75390625,0.18359375,0.09765625}\bm{L_{g}}}\sigma_{g}^{4/3}}{\varepsilon^{10/3}}}}\Bigr{)}$
yes Lip. $\nabla f$ and $\nabla^{2}f$
@@ 2.2 Momentum to the rescue
Besides, its theoretical analysis usually involves assuming that the gradient of $f$ is Lipschitz, an assumption that typically slows down the Cubic Newton method, and which is not required in our work (see also the appendix, where we additionally discuss using the classical momentum in the Cubic Newton).
@@
For illustration, let us assume that the functions $\pi_{\xi}$ are all affine, $\pi_{\xi}({\bm{x}})={\bm{A}}{\bm{x}}+{\bm{b}}_{\xi}$, with the same matrix ${\bm{A}}$ and different bias parts ${\bm{b}}_{\xi}$ (e.g. $\pi_{\xi}$ are the gradients of a quadratic function).
Then we can easily show that
@@
Many combinations of momentum for the gradient and Hessian are possible, but after some investigation, only (IT) for the gradient and (HB) for the Hessian led to an improvement and most importantly to convergence for arbitrary batch sizes for the stochastic Cubic Newton.
@@ 3 Main theoretical analysis
Note about noise Assumptions 2.1,2.1. We will assume we can bound the noise at the initial state ${\bm{x}}_{0}$ slightly differently, and denote $\sigma_{g,0},\sigma_{h,0}$ its corresponding gradient and Hessian noise levels. We do this because, in some cases of our analysis (especially the convex case), the initial noise can give rise to slower terms in our convergence rates, such terms can be easily dealt with by making sure the initial gradient and Hessian estimates are of good quality (smaller variance), which can be done using large batches but only for the initial state. However, we stress that we do not need such a trick and can always guarantee convergence for any batch size even without it.
@@ 3.1 Momentum bounds for Cubic Newton
Lemma 3.1 shows that the average error of (10) can be decomposed into two terms: (variance) term $\alpha^{3/4}\sigma_{g}^{3/2}$ and a (bias) $\frac{\sigma_{g,0}^{3/2}}{\alpha T}+\frac{(1-\alpha)^{3/2}L^{3/2}}{\alpha^{3}}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}r_{t}^{3}$. The (variance) term shows a smaller variance for a value of $\alpha$ that is small, while the (bias) term shows the opposite effect, this means that we should find a sweet spot to balance both terms. Note how the bias can be also decomposed into the initial bias $\frac{\sigma_{g,0}^{3/2}}{\alpha T}$ and the rest as (iterate bias) (iterates are different). The initial bias can be controlled using the initial gradient sample if needed, whereas the iterate bias can be controlled using the term $Mr_{t}^{3}$ that we left in Lemma 2.1 (the main reason why other momentums do not work, in theory, is that they lead to an iterate bias with a power of $r_{t}$ different than $3$).
@@
For example, we can take:
$\begin{array}[]{rcl}\alpha&=&10\sqrt{\frac{L}{M}}\quad\text{and}\quad\beta\;\;=\;\;46\frac{L}{M},\end{array}$ (13)
and to ensure $\alpha,\beta\leq 1$ it suffices to take $M\geq 100L$.
@@ 3.2 Non-convex rates
$\mathbb{E}\mu_{M}({\bm{x}}_{out})=\mathcal{O}\biggl{(}\frac{\sqrt{M}\tilde{F}_{0}}{T}+\frac{L^{3/2}\sigma_{h}^{3}}{M^{3}}+\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}\biggr{)},$
where $\tilde{F}_{0}:=F_{0}+\frac{\sigma_{g,0}^{3/2}}{\sqrt{L}}+\frac{\sigma_{h,0}^{3}}{LM}.$ Note that the quantity $\tilde{F}_{0}$ in Corrolary 3.2 is slightly different from the usual constant $F_{0}$, but their difference is bounded by a constant that depends on the initial gradient and Hessian noise, which means that we can make it as small as possible by for example assuming the first batch is proportional to $T$ in size (e.g. $T/100$); however, we do not have to do this here. By choosing $M$ such that it minimizes the right-hand side in Corrolary 3.2, we get:
$\mathbb{E}\mu_{M}({\bm{x}}_{out})=\mathcal{O}\biggl{(}\frac{\sqrt{L}\tilde{F}_{0}}{T}\\ +\frac{L^{3/14}\tilde{F}_{0}^{6/7}\sigma_{h}^{3/7}}{T^{6/7}}+\frac{L^{3/14}\tilde{F}_{0}^{3/7}\sigma_{g}^{6/7}}{T^{3/7}}\biggr{)}\,.$ (14)
The best old rate in $T$ is $\mathcal{O}(\frac{\sqrt{L}F}{T}+\big{(}\frac{F\sigma_{h}}{T}\big{)}^{3/4}+\sigma_{g}^{3/2})$. Note that we improve both the $\sigma_{h}$ term from $\mathcal{O}(T^{-3/4})$ to $\mathcal{O}(T^{-6/7})$ and most importantly we change the constant term $\mathcal{O}(\sigma_{g}^{3/2})$ to a decreasing term $\mathcal{O}(T^{-3/7})$, thus guaranteeing convergence for any batch size.
@@ 3.3 Extension to the convex case
We observe the same problem as in the non-convex case: the method converges only to a neighborhood of the solution, and it cannot achieve accuracy $\varepsilon\leq\sigma_{g}D$. For the sake of simplicity, we will assume the initial noise constants $\sigma_{g,0},\sigma_{h,0}$ are small enough that we can ignore them (the full, complicated, statement including these terms can be found in the Appendix). {framedtheorem} For any precision $\varepsilon>0$, when $f$ is convex, there exists $M_{\varepsilon}$ such that Algorithm 1 reaches an iterate ${\bm{x}}_{out}^{T}$ with $\mathbb{E}[f({\bm{x}}^{T}_{out})]-f^{\star}\leq\varepsilon$, in a number of iterations at most
$T=\mathcal{O}\Big{(}\log\bigl{(}\frac{F_{0}}{\varepsilon}\bigr{)}+\frac{\sqrt{LD^{3}}}{\sqrt{\varepsilon}}\\ +\frac{D^{7/4}L^{1/4}\tilde{\sigma}_{h}^{1/2}}{\varepsilon^{3/4}}+\frac{D^{5/2}L^{1/4}\sigma_{g}}{\varepsilon^{5/2}}\Big{)}.$
@@ 5 Discussion and Limitations
Other momentum choices. As shown in Subsection 2.2, there are other types of momentum that we tried and for which we were unsuccessful in obtaining improvements, the main reason for this is that we did not have corresponding powers of $r_{t}$ in the iterate bias that let us compensate for the bias using the term $Mr_{t}^{3}$ in Lemma 2.1. we note that this exact problem is present for first-order methods (although it is less complicated since we know $r_{t}$ explicitly in this case) and the fix was to use normalized methods such as in (Cutkosky and Mehta, 2020b, ) to make the size of the step constant. Such a technique is still absent for second-order methods, and were it to be discovered, it would mean the possibility of using the other types of momentum we discussed in this work.
@@ Appendix A Additional results using other momentums
We have also considered other combinations of momentum for the gradient estimation in the Cubic Newton method. We found that they all guarantee convergence to arbitrary accuracy for any batch size. However, they make the Hessian term in the convergence rate slower, and only in the case of the MVR momentum are we able to show an improved rate for the gradient term.
@@ C.3 Nonconvex convergence
Choosing $\alpha,\beta\in(0,1]$ such that:
$\frac{4\sqrt{3}L^{3/2}}{\alpha^{3}}+\frac{73}{M^{3/2}}\frac{9L^{3}}{\beta^{3}}-\frac{M^{3/2}}{72}\leq 0$ (24)
For example, we can take: $\alpha=10\sqrt{\frac{L}{M}}$ and $\beta=46\frac{L}{M}$, to ensure that both are smaller than one it suffices to take $M\geq 100L$.
@@
$\frac{1}{1008}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mu_{M}({\bm{x}}_{t+1})$ $\leq\frac{\sqrt{M}\tilde{F_{0}}(M)}{T}+40\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}+\frac{73(46)^{3/2}L^{3/2}}{M^{3}}\tilde{\sigma}_{h}^{3}(\beta),$
where $\tilde{F_{0}}(M):=F_{0}+1.2\frac{\sigma_{g,0}^{3/2}}{\sqrt{L}}+\frac{16}{ML}\sigma_{h,0}^{3}$ and $\tilde{\sigma}_{h}^{3}(\beta)=8(er)^{3/2}\sigma_{h}^{3}+16(46)^{3/2}(er)^{3}\frac{L^{3/2}}{M^{3/2}}\delta_{h}^{3}$ (remember that $r=2\log(d)$).
@@
We clearly have a better dependence on the regularization parameter $M$ in both the gradient and Hessian noise terms.
@@
$\frac{1}{1008}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mu_{M}({\bm{x}}_{t+1})$ $\leq\frac{\sqrt{M}\tilde{F_{0}}(M)}{T}+40\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}+\frac{73(46)^{3/2}L^{3/2}}{M^{3}}\tilde{\sigma}_{h}^{3}(\beta).$
@@
$M=\max\Big{(}100L,\frac{L^{3/7}\sigma_{g}^{12/7}}{\tilde{F_{0}}^{8/7}}T^{8/7},\frac{L^{3/7}\tilde{\sigma}_{h}(T)^{3/7}}{\tilde{F_{0}}^{2/7}}T^{2/7}\Big{)}$
@@
$\frac{1}{1008}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mu_{M}({\bm{x}}_{t+1})$ $\leq\frac{\sqrt{L}\tilde{F_{0}}}{T}+40\frac{L^{3/14}\tilde{F_{0}}^{3/7}\sigma_{g}^{6/7}}{T^{3/7}}+73(46)^{3/2}\frac{L^{3/14}\tilde{F_{0}}^{6/7}\tilde{\sigma}_{h}(T)^{3/7}}{T^{6/7}}.$

### v2

@@ 2.2 Momentum to the rescue
Besides, its theoretical analysis usually involves assuming the gradient of $f$ is Lipschitz, an assumption that typically slows down the Cubic Newton method, and which is not required in our work (see also the appendix, where we additionally discuss using the classical momentum in the Cubic Newton).
@@
For illustration, let us assume that the functions $\pi_{\xi}$ are all affine, $\pi_{\xi}({\bm{x}})={\bm{A}}{\bm{x}}+{\bm{b}}_{\xi}$, with the same matrix ${\bm{A}}$ and different bias parts ${\bm{b}}_{\xi}$ (e.g. $\pi_{\xi}$ are the gradients of a quadratic function). Then we can easily show that
@@
Many combinations of momentum for the gradient and Hessian are possible but we will only consider in the main paper, (IT) for the gradient and (HB) for the Hessian as this is the only combination that led to an improvement and most importantly to convergence for arbitrary batch sizes for the stochastic Cubic Newton without needing any additional assumption. We consider other combinations in the appendix.
@@ 3 Main theoretical analysis
Note about noise Assumptions 2.1,2.1. We will assume we can bound the noise at the initial state ${\bm{x}}_{0}$ slightly differently, and denote $\sigma_{g,0}\leq\sigma_{g},\sigma_{h,0}\leq\sigma_{h}$ its corresponding gradient and Hessian noise levels. We do this because, in some cases of our analysis (especially the convex case), the initial noise can give rise to slower terms in our convergence rates, such terms can be easily dealt with by making sure the initial gradient and Hessian estimates are of good quality (smaller variance), which can be done using large batches but only for the initial state. However, we stress that we do not need such a trick and can always guarantee convergence for any batch size even without it.
@@ 3.1 Momentum bounds for Cubic Newton
Lemma 3.1 shows that the average error of (10) can be decomposed into two terms: (variance) term $\alpha^{3/4}\sigma_{g}^{3/2}$ and a (bias) $\frac{\sigma_{g,0}^{3/2}}{\alpha T}+\frac{(1-\alpha)^{3/2}L^{3/2}}{\alpha^{3}}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}r_{t}^{3}$. The (variance) term shows a smaller variance for a value of $\alpha$ that is small, while the (bias) term shows the opposite effect, this means that we should find a sweet spot to balance both terms. Note how the bias can also be decomposed into the (initial bias) $\frac{\sigma_{g,0}^{3/2}}{\alpha T}$ and the rest as (iterate bias) (iterates are different). The initial bias can be controlled using the initial gradient sample if needed. In contrast, the iterate bias can be controlled using the term $Mr_{t}^{3}$ that we left in Lemma 2.1 (the main reason why other momentums do not work, in theory, is that they lead to an iterate bias with a power of $r_{t}$ different than $3$).
@@
To choose the momentum parameters $\alpha,\beta$, we proceed as follows:
@@
To satisfy (12), it suffices to take:
$\begin{array}[]{rcl}\alpha&\geq&10\sqrt{\frac{L}{M}}\quad\text{and}\quad\beta\;\;\geq\;\;46\frac{L}{M},\end{array}$ (13)
Controlling the initial bias terms. To accomplish this, we further choose
$\begin{array}[]{rcl}\alpha&\in&\underset{\alpha\in(0,1]}{{\rm arg\,min}}\;\;\alpha^{3/4}\sigma_{g}^{3/2}+\frac{\sigma_{g,0}^{3/2}}{\alpha T}=\frac{a_{g}^{6/7}}{T^{4/7}}\\ &\text{and}&\\ \beta&\in&\underset{\beta\in(0,1]}{{\rm arg\,min}}\;\;\beta^{3/2}\sigma_{h}^{3}+\frac{1}{\beta T}\sigma_{h,0}^{3}=\frac{a_{h}^{6/5}}{T^{2/5}},\end{array}$ (14)
where, for simplicity, we denoted $a_{g}=\frac{\sigma_{g,0}}{\sigma_{g}}\leq 1$ and $a_{h}=\frac{\sigma_{h,0}}{\sigma_{h}}\leq 1$.
All in all, based on (13) and (14), we choose
$\begin{array}[]{rcl}\alpha=\max\Big{(}\frac{a_{g}^{6/7}}{T^{4/7}},10\sqrt{\frac{L}{M}}\Big{)}\,,\,\beta=\max\Big{(}\frac{a_{h}^{6/5}}{T^{2/5}},46\frac{L}{M}\Big{)}\,.\end{array}$ (15)
To ensure $\alpha,\beta\leq 1$ it suffices to take $M\geq 100L$.
@@ 3.2 Non-convex rates
$\mathbb{E}\mu_{M}({\bm{x}}_{out})=\mathcal{O}\biggl{(}\frac{\sqrt{M}F_{0}}{T}+\frac{L^{3/2}\sigma_{h}^{3}}{M^{3}}+\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}\\ +\frac{a_{g}^{9/14}\sigma_{g}^{3/2}}{T^{3/7}}+\frac{a_{h}^{9/5}\sigma_{h}^{3}}{M^{3/2}T^{3/5}}\biggr{)}.$
The last two terms represent the effect of the initial gradient and Hessian noise on convergence and we can see that they decay with $T$.
By choosing $M$ such that it minimizes the right-hand side in Corollary 3.2, we get:
$\mathbb{E}\mu_{M}({\bm{x}}_{out})=\mathcal{O}\biggl{(}\frac{\sqrt{L}F_{0}}{T}\\ +\frac{L^{3/14}F_{0}^{6/7}\sigma_{h}^{3/7}}{T^{6/7}}+\frac{L^{3/14}F_{0}^{3/7}\sigma_{g}^{6/7}}{T^{3/7}}\\ +\frac{a_{g}^{9/14}\sigma_{g}^{3/2}}{T^{3/7}}+\frac{a_{h}^{9/20}\sigma_{h}^{3/4}F_{0}^{1/2}}{T^{9/10}}\biggr{)}\,.$ (16)
The best old rate in $T$ is $\mathcal{O}(\frac{\sqrt{L}F_{0}}{T}+\big{(}\frac{F_{0}\sigma_{h}}{T}\big{)}^{3/4}+\sigma_{g}^{3/2})$. Note that we improve both the $\sigma_{h}$ term from $\mathcal{O}(T^{-3/4})$ to $\mathcal{O}(T^{-6/7})$ and, most importantly, we change the constant term $\mathcal{O}(\sigma_{g}^{3/2})$ to a decreasing term $\mathcal{O}(T^{-3/7})$, thus guaranteeing convergence for any batch size.
@@ 3.3 Extension to the convex case
We observe the same problem as in the non-convex case: the method converges only to a neighborhood of the solution, and it cannot achieve accuracy $\varepsilon\leq\sigma_{g}D$.{framedtheorem} For $\alpha,\beta$ defined in (15) and $M\geq 100L$, when the objective $f$ is convex, the iterates of Algorithm 1 satisfy:
$\min_{0\leq t\leq T-1}F_{t}=\mathcal{O}\Big{(}\frac{MD^{3}}{T^{2}}+\frac{LD\tilde{\sigma}_{h}^{2}}{M^{2}}+\frac{L^{1/4}}{M^{1/4}}D\sigma_{g}\\ +F_{0}e^{-T/3}+\frac{a_{g}^{3/7}D\sigma_{g}}{T^{2/7}}+\frac{a_{h}^{6/5}D\sigma_{h}^{2}}{MT^{2/5}}\Big{)},$
where we define $F_{t}=\mathbb{E}[f({\bm{x}}_{t})]-f^{\star}$.
Choosing $M$ such that it minimizes the right-hand side, we get:
$\min_{0\leq t\leq T-1}F_{t}\\ =\mathcal{O}\Big{(}\frac{LD^{3}}{T^{2}}+\frac{L^{1/3}D^{7/3}\tilde{\sigma}_{h}^{2/3}}{T^{4/3}}+\frac{L^{1/5}D^{7/5}\sigma_{g}^{4/5}}{T^{2/5}}\\ +F_{0}e^{-T/3}+\frac{a_{h}^{3/5}D^{2}\sigma_{h}}{T^{6/5}}+\frac{a_{g}^{3/7}D\sigma_{g}}{T^{2/7}}\Big{)}.$
The last three terms in Theorem 3.3 correspond to the effect of the initialization and the initial gradient and Hessian noises. The last two terms can be made even smaller by choosing a larger initial batch (i.e. reducing $a_{g},a_{h}$).
@@
By ignoring the last three terms, we get the following iteration complexity: $\mathcal{O}\Big{(}(\frac{LD^{3}}{\varepsilon})^{1/3}+\frac{L^{1/4}D^{7/4}\tilde{\sigma}_{h}^{1/2}}{\varepsilon^{3/4}}+\frac{L^{1/2}D^{7/2}\sigma_{g}^{2}}{\varepsilon^{5/2}}\Big{)}$.
@@ 5 Discussion and Limitations
Practical implementation. We did not provide a specific method for solving the cubic sub-problem in Algorithm 1. A naive implementation would require access to the entire Hessian matrix, making it impractical for high-dimensional problems. However, we can solve this problem using only Hessian-vector products, which are typically more efficient for various problem classes, such as neural networks with backpropagation. It is important to note that incorporating Hessian momentum may complicate this process. Therefore, we leave the exploration of how to effectively combine Hessian-vector products and momentum for future work.
Other momentum choices. As shown in Subsection 2.2, there are other types of momentum that we tried in the supplementary material. We were able to improve the gradient noise term from $\mathcal{O}(\frac{1}{T^{3/7}})$ in (16) to $\mathcal{O}(\frac{1}{T^{9/20}})$ using the MVR momentum from (Cutkosky and Orabona,, 2020); this improvement comes at the cost of needing to assume the gradient of the objective is $L_{g}$-Lipschitz, but this constant can be made as big as we want by adding quadratic functions to our objective. Even when ignoring this problem with the constant $L_{g}$, the corresponding iteration complexity is only $\mathcal{O}(\frac{1}{\varepsilon^{30/9}})$, which is worse than $\mathcal{O}(\frac{1}{\varepsilon^{3}})$ obtained using the corresponding momentum in first-order methods (Cutkosky and Orabona,, 2020). One of the reasons we think this is the case is that we did not have corresponding powers of $r_{t}$ in the iterate bias that let us compensate for the bias using the term $Mr_{t}^{3}$ in Lemma 2.1. We note that this exact problem is present for first-order methods (although it is less complicated since we know $r_{t}$ explicitly in this case) and the fix was to use normalized methods such as in (Cutkosky and Mehta, 2020b, ) to make the size of the step constant. Such a technique is still absent for second-order methods, and were it to be discovered, it would mean the possibility of improving the other types of momentum we discussed in this work.
@@ 6 Conclusion
Acknowledgments
This work was supported by the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number 22.00133.
@@ Appendix A Additional results using other momentums
We have also considered other combinations of momentum for the gradient estimation in the Cubic Newton method. We found out that they all guarantee convergence to arbitrary accuracy for any batch size. However, they make the Hessian term in the convergence rate slower, and only in the case of the MVR momentum we are able to show an improved rate for the gradient term.
@@ C.3 Nonconvex convergence
We choose $\alpha,\beta\in(0,1]$ such that:
$\frac{4\sqrt{3}L^{3/2}}{\alpha^{3}}+\frac{73}{M^{3/2}}\frac{9L^{3}}{\beta^{3}}-\frac{M^{3/2}}{72}\leq 0\;,$ (26)
and to deal with initial gradient and Hessian noises, we also take:
$\alpha\in{\rm arg\,min}\frac{12}{\alpha T}\sigma_{g,0}^{3/2}+4\sqrt{3}\alpha^{3/4}\sigma_{g}^{3/2}\approx\frac{a_{g}^{6/7}}{T^{4/7}}$ and $\beta\in{\rm arg\,min}\frac{730}{M^{3/2}\beta T}\sigma_{h,0}^{3}+\frac{73}{M^{3/2}}\beta^{3/2}\sigma_{h}^{3}\approx\frac{a_{h}^{6/5}}{T^{2/5}}$.
Where, for simplicity, we denoted $a_{g}=\frac{\sigma_{g,0}}{\sigma_{g}}\leq 1$ and $a_{h}=\frac{\sigma_{h,0}}{\sigma_{h}}\leq 1$.
All in all, we choose
$\begin{array}[]{rcl}\alpha=\max\Big{(}\frac{a_{g}^{6/7}}{T^{4/7}},10\sqrt{\frac{L}{M}}\Big{)}\,,\,\beta=\max\Big{(}\frac{a_{h}^{6/5}}{T^{2/5}},46\frac{L}{M}\Big{)}\,.\end{array}$ (27)
To ensure $\alpha,\beta\leq 1$ it suffices to take $M\geq 100L$.
@@
$\frac{1}{1008}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mu_{M}({\bm{x}}_{t+1})$ $\leq\frac{\sqrt{M}F_{0}}{T}+40\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}+\frac{73(46)^{3/2}L^{3/2}}{M^{3}}\tilde{\sigma}_{h}^{3}(\beta)+20\frac{a_{g}^{9/14}\sigma_{g}^{3/2}}{T^{3/7}}+803\frac{a_{h}^{9/5}\sigma_{h}^{3}}{M^{3/2}T^{3/5}},$
where $\tilde{\sigma}_{h}^{3}(\beta)=8(er)^{3/2}\sigma_{h}^{3}+16(46)^{3/2}(er)^{3}\frac{L^{3/2}}{M^{3/2}}\delta_{h}^{3}$ (remember that $r=2\log(d)$).
@@
We clearly depend better on the regularization parameter $M$ in both the gradient and Hessian noise terms.
@@
$\frac{1}{1008}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mu_{M}({\bm{x}}_{t+1})$ $\leq\frac{\sqrt{M}F_{0}}{T}+40\frac{L^{3/8}}{M^{3/8}}\sigma_{g}^{3/2}+\frac{73(46)^{3/2}L^{3/2}}{M^{3}}\tilde{\sigma}_{h}^{3}(\beta)+20\frac{a_{g}^{9/14}\sigma_{g}^{3/2}}{T^{3/7}}+803\frac{a_{h}^{9/5}\sigma_{h}^{3}}{M^{3/2}T^{3/5}}.$
@@
$M=\max\Big{(}100L,\frac{L^{3/7}\sigma_{g}^{12/7}}{F_{0}^{8/7}}T^{8/7},\frac{L^{3/7}\tilde{\sigma}_{h}(T)^{3/7}}{F_{0}^{2/7}}T^{2/7},\frac{a_{h}^{9/10}\sigma_{h}^{3/2}T^{1/5}}{\sqrt{F_{0}}}\Big{)}$
@@
$\frac{1}{1008}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mu_{M}({\bm{x}}_{t+1})$ $\leq\frac{\sqrt{L}F_{0}}{T}+\frac{40L^{3/14}F_{0}^{3/7}\sigma_{g}^{6/7}+20a_{g}^{9/14}\sigma_{g}^{3/2}}{T^{3/7}}+73(46)^{3/2}\frac{L^{3/14}F_{0}^{6/7}\tilde{\sigma}_{h}(T)^{3/7}}{T^{6/7}}+804\frac{a_{h}^{9/20}\sigma_{h}^{3/4}F_{0}^{1/2}}{T^{9/10}}.$